OpenStudy (anonymous):

g(t)= -t^2 - 2t h(t)= 3t-3 Find (g-h)(t+2) i just need to see if my answer is correct.

OpenStudy (anonymous):

@jim_thompson5910 thia one is really challenging.

OpenStudy (anonymous):

this*

jimthompson5910 (jim_thompson5910):

hint: (g-h)(t+2) is the same as g(t+2) - h(t+2)

OpenStudy (anonymous):

I know but the algebra sometimes trips me up. my final answer came out to t^2-9t-9 but i have a feeling i messed up somewhere along the line

jimthompson5910 (jim_thompson5910):

that's pretty close, but not quite correct

OpenStudy (anonymous):

ahh darn

jimthompson5910 (jim_thompson5910):

g(t)= -t^2 - 2t g(t+2)= -(t+2)^2 - 2(t+2) g(t+2)= -(t^2+4t+4) - 2(t+2) g(t+2)= -t^2-4t-4 - 2t-4 g(t+2)= -t^2-6t-8 --------------------------------- h(t)= 3t-3 h(t+2)= 3(t+2)-3 h(t+2)= 3t+6-3 h(t+2)= 3t+3 =========================== (g-h)(t+2) g(t+2) - h(t+2) (-t^2-6t-8) - (3t+3) I'll let you finish

OpenStudy (anonymous):

whew is it -t^2 -9t-5

jimthompson5910 (jim_thompson5910):

close, but - (3t+3) becomes - 3t-3

jimthompson5910 (jim_thompson5910):

NOT -3t + 3

OpenStudy (anonymous):

oh so is it -t^2 -3t -5 ?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

(-t^2-6t-8) - (3t+3) -t^2-6t-8 - 3t-3

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

let's try again. -t^2-9t-11 or -t^2-3t-11

jimthompson5910 (jim_thompson5910):

-6 - 3 = ??

OpenStudy (anonymous):

-9

jimthompson5910 (jim_thompson5910):

so -6t - 3t = -9t

jimthompson5910 (jim_thompson5910):

it's the first one

OpenStudy (anonymous):

thanks so so so so so much for bearing with me and not stabbing your eyes out.

jimthompson5910 (jim_thompson5910):

lol i like my eyes, so I wouldn't do that, but yw